3(x+1)=3(x+7)
Choose the solution
No solution
all real numbers

Answer by Mathtut(3670)

3(x+1)=3(x+7)
:
distribute right and left sides
:
3x+3=3x+21
:
subtract 3x from both sides
:
3=21......which is an untrue statement .....therefore NO solutions.